# Meridional Circulation Streamlined

### v1

@@ Abstract
Time-dependent meridional circulation and differential rotation in radiative zones are central open issues in stellar evolution theory. We streamline this challenging problem using the ‘downward control principle’ of atmospheric science, under a geostrophic f-plane approximation. New insights emerge from this simplified approach, using pressure as the vertical coordinate. We recover the known stellar physics result that the steady-state meridional circulation decays on the length scale $(N/2\Omega\times\sqrt{Pr})H$, assuming molecular viscosity is the dominant drag mechanism. Prior to steady-state, the meridional circulation and the zonal wind (= differential rotation) spread together via radiative diffusion, under thermal wind balance. The corresponding (4th-order) hyperdiffusion process is reasonably well approximated by regular (2nd-order) diffusion on scales of order a pressure scale-height. We derive an inhomogeneous diffusion (equiv. advection-diffusion) equation for the zonal flow which admits closed-form time-dependent solutions in a finite depth domain, allowing rapid prototyping of the meridional circulation pattern. In the weak drag limit, we find that the time to rotational steady-state can be longer than the Eddington-Sweet time and be instead determined by the longer drag time. Unless strong enough drag operates, the internal rotation of main-sequence stars may thus never reach steady-state. Streamlined meridional circulation solutions provide leading-order internal rotation profiles for studying the role of fluid/MHD instabilities (or waves) in redistributing angular momentum in the radiative zones of stars. Despite clear geometrical limitations and simplifying assumptions, one might expect our thin-layer geostrophic approach to offer qualitatively useful results to understand deep meridional circulation in stars.
@@ 1 Introduction
Helioseismology (Christensen-Dalsgaard, 2002) shows that the outer convective envelope of the Sun(0.7 - 1 $R_{\odot}$), is rotating differentially (Schou et al., 1998), while the radiative zone (0.2 - 0.7 $R_{\odot}$) is in approximate solid body rotation (Brown et al., 1989; Goode et al., 1991). The differential rotation in the convective zone is attributed to different processes like varying Reynolds stresses arising from turbulent convection (Rüdiger, 2022) or torques induced by solar winds (Modisette, 1967; Weber and Davis Jr, 1967). Therefore, the radiative zone may be conceived as being torqued by the convection zone from above, leading to the generation of meridional circulation via. a mechanism called gyroscopic pumping (Hughes et al., 2007; Garaud and Arreguin, 2009; Garaud and Bodenheimer, 2010).
@@
It turns out that there is a closely analogous problem in atmospheric science: the downward control principle (Haynes et al., 1991) proposed to describe the effect of torques due to waves that are dissipated in the upper atmosphere (stratosphere) of the Earth. As the name suggests, torques on the low-density upper layers can exert control on the otherwise inert and dense deeper layers. The process can be understood as follows: alternating zonal flows pumped in the upper layer get deflected by Coriolis forces along the longitudinal direction similar to Ekman flows (Liu et al., 2021) in the ocean. This results in meridionally periodic regions of mass convergence and divergence at the forced level, thereby generating barotropic pressure gradients on all constant-height surfaces “throughout the depth of the fluid” (Showman et al., 2006). The pressure gradient causes a return flow in the unforced deeper layers, completing the meridional circulation loop. This return flow, in turn, is again deflected by Coriolis forces, generating a deep zonal flow in the same direction as the top-forced wind. Thus, meridional circulation and zonal flow (differential rotation) penetrate jointly through the depth of the rotating stratified fluid.
@@
Meridional circulation of Sun-like stars has seen several analytical (Eddington, 1925; Sweet, 1950; Spiegel and Zahn, 1992; Garaud and Brummell, 2008) and numerical (Clement, 1993; Talon et al., 2003; Garaud and Arreguin, 2009) treatments under the Boussinesq (Rieutord, 2006) or the Anelastic (Garaud, 2002) approximations inspired from geophysical fluid dynamics. Despite these efforts, the process of penetration arguably lacks a simple, intuitive explanation in the stellar astrophysics literature. Using the aforementioned analogy from atmospheric sciences, we re-consider the problem of meridional circulation in stars and the related issue of the solar tachocline. We argue that the greater physical intuition gained from the atmospheric science perspective outweighs its geometric/simplifying assumptions, making way for a better qualitative understanding of the process. Additionally, we highlight the importance of time-dependent solutions (Elliott, 1997; Wood and Brummell, 2012, 2018) and extend its applications to the context of stellar evolution codes.
@@
We adopt the primitive flimit-$f-$plane meteorological framework of Showman et al. (2006), which investigates the depth of penetration of zonal jets and their associated meridional circulation into the molecular envelope of Jupiter. In §2, we discuss the assumptions and derive an expression for the steady-state meridional streamfuction, thereby confirming that it decays on the length scale $(N/2\Omega\times\sqrt{Pr})H$ (Spiegel and Zahn, 1992; Garaud, 2002). In §3, we derive a single 1D advection-diffusion equation for the zonal velocity in log-pressure coordinates and explain both steady-state and time-dependent analytical solutions, and compare them with Dedalus2 numerical solutions. Finally, we conclude in §4.
@@ 2 The meridional streamfunction
Figure 1a shows the internal structure of a solar-type star with an outer convective zone, an inner radiative zone, and the core at the center. The zonal, meridional, and vertical coordinates are $\phi,\theta$, and $z$, respectively.
@@ 2.1 Assumptions
  • f-plane (see figure 1a): We use Cartesian coordinates rather than spherical. The f-plane approximation uses a fixed value of the Coriolis parameter $f(=2\Omega\sin\theta)$ set to correspond to mid-latitudes ($\theta=\theta_{0}$). Here $\Omega$ is the background rotation rate.
@@
  • Geostrophic balance: The Rossby number $Ro\sim\Delta\Omega/2\Omega$ of the flow is small. Here, $\Delta\Omega$ is the deviation of the latitudinal differential rotation from the background $\Omega$. Inertial forces are therefore dominated by rotation $\Omega$.
@@ 2.2 The steady-state streamfunction
The independent variables representing the vertical and the meridional coordinates are pressure $p$ and distance $y$, respectively. Time is denoted by $t$. The zonal-mean momentum, thermal wind, mass continuity and thermodynamic energy equations for a stratified hydrostatic atmosphere with momentum forcing are given by, respectively,
@@
where $u$, $v$, and $w(\equiv\partial p/\partial t)$ are the zonal, meridional, and vertical wind speeds in the rotating frame of reference, $f(=2\Omega\sin\theta)$ is the Coriolis frequency parameter, $\theta$ is the latitude and $R$ is the ideal gas constant. The momentum forcing relaxes $u$ towards $u_{F}$, and the thermal forcing relaxes $T$ towards $T^{\prime}_{E}$. The corresponding frictional drag ($\tau_{F}$) and radiative ($\tau_{R}$) relaxation times approximate the usual operators for viscous and heat diffusion. The Brunt frequency $N$ for the stratified radiative zone with a temperature distribution $T=T(p)+T^{\prime}$, where $T^{\prime}$ is a small perturbation from the reference state $T$, is given by
@@
where $C_{P}$ is the specific heat of the constituent gas at constant pressure or along isobars, $H=RT/g$ is the isothermal scale height, and $g$ is the acceleration due to gravity. The advection terms in the momentum equation can be ignored in first approximation owing to the low Rossby number $Ro$ of the flow resulting from geostrophic balance (Showman et al., 2006). We thus have four equations (1)-(4) and four variables $u$, $v$, $w$ and $T^{\prime}$.
In this work, we consider thermal forcing to be zero i.e. $T^{\prime}_{E}=0$ to focus on the role of momentum forcing for meridional circulation. On eliminating $u$ and $T^{\prime}$ from (1)-(4) and substituting $v$ and $w$ with their corresponding expressions in terms of the meridional streamfunction $\psi$ ($v=-\partial\psi/\partial p$ and $w=\partial\psi/\partial y$), we obtain
@@
Our domain of interest is divided into two parts: the top layer corresponds to the near-convection zone where momentum forcing is applied (RHS of equation 1), while the deeper, stably stratified layer below is passive and only driven by the circulation in the top forced layer. The forcing velocity $u_{F}$, relaxation timescales $\tau_{R}$ and $\tau_{F}$, and the Rossby deformation radius $L_{D}(=NH/f)$ are piecewise constant along the vertical coordinate $p$, demarcating the two layers described. In the passive layer, where $u_{F}=0$, equation (6) reduces to,
$L_{D}^{2}\frac{\partial^{2}\psi}{\partial y^{2}}+p^{2}\frac{\tau_{F}}{\tau_{R}}\frac{\partial^{2}\psi}{\partial p^{2}}=0$ (7)
This equation, together with appropriate boundary conditions, gives the steady-state meridional streamfunction in a rotating hydrostatically and geostrophically balanced fluid such as the radiative zone of a Sun-like star. We see that $\psi$ admits a separable solution of the form $\psi(p,y)=\psi_{p}(p)\psi_{y}(y)$, with $\psi_{y}(y)=\cos(ly)$ i.e. periodic in latitude, where $l(=1/L_{y})$ can represent the latitudinal wavenumber of the shallow jets forced in the top layer. Substituting the above in equation (7) we obtain an ordinary differential equation for the vertical component of the streamfunction $\psi_{p}$ as
$p^{2}\frac{\partial^{2}\psi_{p}}{\partial p^{2}}=\bar{\omega}^{2}\psi_{p}$ (8)
where $\bar{\omega}^{2}=L_{D}^{2}l^{2}\tau_{R}/\tau_{F}$. It admits the polynomial solution for $\psi$:
@@
where, $n_{1}$ and $n_{2}$ are roots of the equation $n(n-1)=\bar{\omega}^{2}$. Constants $A$ and $B$ may be determined from interface and boundary conditions. The vertical component of the steady-state streamfunction, which determines the steady-state depth of the meridional circulation, is therefore strictly dependent on $\bar{\omega}$ which is the product of the Burger number (Bu = $L_{D}^{2}l^{2}$) and the Prandlt number (Pr = $\tau_{R}/\tau_{F}$) as has already been established in literature e.g. Spiegel and Zahn (1992); Garaud (2002); Gilman and Miesch (2004); Garaud and Brummell (2008). We now use this setup and formalism to explore the time-dependent behavior of the meridional circulation in a top-forced radiative zone.
@@ 3 Time-dependent zonal velocity equation
$f\frac{\partial\bar{u}}{\partial\ln p}=lR\bar{T}$ (12)
@@
where, $L_{D}=(NH)/f$ is the Rossby deformation radius, the log-pressure vertical coordinate $z=\ln p$ and $k=1/(L_{D}^{2}l^{2}\tau_{R})$. It can be rewritten as a linear advection-diffusion equation for the evolution of the zonal wind (RHS of equation 15). Accounting for the simplification of the radiative diffusion operator as Newtonian linear relaxation, equation (15) shares close similarity with the hyperdiffusion equation obtained for the radial profile of differential rotation by Spiegel and Zahn (1992). In its most general form, equations (15) may have $L_{D},\tau_{R},\tau_{F}$, and $k$ that dependent on $p$. However, it is possible to find useful approximate analytical solutions to equation (15) if the above parameters are constant. We apply the Dirichlet boundary condition $\bar{u}(0,t)=u_{F}$ at the top of the passive domain as an idealized form of zonal forcing. The bottom of the domain has a free slip (no shear) boundary, and hence a Neumann boundary condition given by $u_{z}(L,t)=0$. Under these conditions, one can use solution C3 in Van Genuchten (1982) for the steady-state,
@@
$\bar{u}(z,t)=\frac{u_{F}B(z,t)}{2C(z)};m^{\pm}=\exp{\left(-\frac{k\pm j}{2k}\right)};n^{\pm}=\frac{z\pm jt}{2\sqrt{+kt}};\\ B(z,t)=m^{-}\operatorname{erfc}(n^{-})+m^{+}\operatorname{erfc}(n^{+})+\exp{\left(\frac{jL}{k}\right)}\times\\ \left[\frac{j+k}{j-k}m^{+}\operatorname{erfc}{\left(\frac{L}{\sqrt{+kt}}-n^{+}\right)}+\frac{j-k}{j+k}m^{-}\operatorname{erfc}{\left(\frac{L}{\sqrt{+kt}}-n^{-}\right)}\right]\\ \frac{-k\tau_{F}}{2}\exp{\left(-L-\frac{t}{\tau_{F}}\right)}\operatorname{erfc}{\left(\frac{2L-z-kt}{2\sqrt{+kt}}\right)}$ (17)

### v2

@@ Abstract
Time-dependent meridional circulation and differential rotation in radiative zones are central open issues in stellar evolution theory. We streamline this challenging problem using the ‘downward control principle’ of atmospheric science, under a geostrophic f-plane approximation. We recover the known stellar physics result that the steady-state meridional circulation decays on the length scale proportional to $N/f\times\sqrt{Pr}$, assuming molecular viscosity is the dominant drag mechanism. Prior to steady-state, the meridional circulation and the zonal wind (= differential rotation) spread together via radiative diffusion, under thermal wind balance. The corresponding (4th-order) hyperdiffusion process is reasonably well approximated by regular (2nd-order) diffusion on scales of order a pressure scale-height. We derive an inhomogeneous diffusion (equiv. advection-diffusion) equation for the zonal flow which admits closed-form time-dependent solutions in a finite depth domain, allowing for rapid prototyping of differential rotation profiles. In the weak drag limit, we find that the time to rotational steady-state can be longer than the Eddington-Sweet time and be instead determined by the longer drag time. Unless strong enough drag operates, the internal rotation of main-sequence stars may thus never reach steady-state. Streamlined meridional circulation solutions provide leading-order internal rotation profiles for studying the role of fluid/MHD instabilities (or waves) in redistributing angular momentum in the radiative zones of stars. Despite clear geometrical limitations and simplifying assumptions, one might expect our thin-layer geostrophic approach to offer qualitatively useful results to understand deep meridional circulation in stars.
@@ 1 Introduction
Helioseismology (Christensen-Dalsgaard, 2002) shows that the outer convective envelope of the Sun(0.7 - 1 $\rm{R}_{\odot}$), is rotating differentially (Schou et al., 1998), while the radiative zone (0.2 - 0.7 $\rm{R}_{\odot}$) is in approximate solid body rotation (Brown et al., 1989; Goode et al., 1991). The differential rotation in the convective zone is attributed to different processes like varying Reynolds stresses arising from turbulent convection (Rüdiger, 2022) or torques induced by solar winds (Modisette, 1967; Weber and Davis Jr, 1967). Therefore, the radiative zone may be conceived as being torqued by the convection zone from above, leading to the generation of meridional circulation via. a mechanism called gyroscopic pumping (Hughes et al., 2007; Garaud and Arreguin, 2009; Garaud and Bodenheimer, 2010).
@@
It turns out that there is a closely analogous problem in atmospheric science: the downward control principle (Haynes et al., 1991) proposed to describe the effect of torques due to waves that are dissipated in the upper atmosphere (stratosphere) of the Earth. As the name suggests, torques on the low-density upper layers can exert ‘control’ on the otherwise inert and dense deeper layers. The process can be understood as follows: alternating zonal flows pumped in the upper layer get deflected by Coriolis forces along the longitudinal direction similar to Ekman flows (Liu et al., 2021) in the ocean. This results in meridionally periodic regions of mass convergence and divergence at the forced level, thereby generating barotropic pressure gradients on all constant-height surfaces “throughout the depth of the fluid” (Showman et al., 2006). The pressure gradient causes a return flow in the unforced deeper layers, completing the meridional circulation loop. This return flow, in turn, is again deflected by Coriolis forces, generating a deep zonal flow in the same direction as the top-forced wind. Thus, meridional circulation and zonal flow (differential rotation) penetrate jointly through the depth of the rotating stratified fluid.
@@
Meridional circulation of Sun-like stars has seen several analytical (Eddington, 1925; Sweet, 1950; Spiegel and Zahn, 1992; Garaud and Brummell, 2008) and numerical (Clement, 1993; Talon et al., 2003; Garaud and Arreguin, 2009) treatments under the Boussinesq (Rieutord, 2006) or the Anelastic (Garaud, 2002) approximations inspired from geophysical fluid dynamics. Despite these efforts, the process of penetration arguably lacks a simple, intuitive explanation in the stellar astrophysics literature. Using the aforementioned analogy from atmospheric sciences, we re-consider the problem of meridional circulation in stars. We argue that the greater physical intuition gained from the atmospheric science perspective outweighs its geometric/simplifying assumptions, making way for a better qualitative understanding of the process.
While most previous studies discuss the steady-state meridional circulation (Garaud and Brummell, 2008; Garaud and Arreguin, 2009), some that focus on the unsteady problem (Elliott, 1997; Wood and Brummell, 2012, 2018) resort to numerical simulations making them computationally expensive. In this work, a first set of analytical solutions is presented for the time-dependent problem, allowing for convenient reproducibility. Such a solution is important as stars are low Prandtl number environments and hence may not reach rotational steady-state in their main-sequence lifetimes. In the context of stellar evolution codes like MESA, the time-dependent analytical solution will serve to accurately transport angular momentum and chemicals through meridional circulation. Additionally, differential rotation profiles obtained may be probed for magnetic/fluid instabilities/waves, thereby allowing the calculation of appropriate diffusion coefficients that parameterize transport beyond what is currently implemented in stellar models.
@@
We adopt the primitive flimit-$f-$plane meteorological framework of Showman et al. (2006), which investigates the depth of penetration of zonal jets and their associated meridional circulation into the molecular envelope of Jupiter. In §2, we discuss the assumptions and derive an expression for the steady-state meridional streamfuction, thereby confirming that it decays on a length scale proportional to $N/f\times\sqrt{Pr}$, a product of buoyancy and rotation frequency ratio and the Prandtl number (Spiegel and Zahn, 1992; Garaud, 2002). In §3, we derive a single 1D advection-diffusion equation for the zonal velocity in log-pressure coordinates and explain both steady-state and time-dependent analytical solutions, comparing them with Dedalus2 numerical solutions in §4. Finally, we conclude in §5.
@@ 2 The meridional streamfunction
Figure 1a shows the internal structure of a solar-type star with an outer convective zone, an inner radiative zone, and the core at the center. The zonal, meridional, and vertical coordinates are $\phi,\theta$, and $z$, respectively. The convective zone applies a zonal torque on the underlying radiative zone represented as prograde and retrograde winds in figure 1b.
@@ 2.1 Assumptions
  • f-plane (see figure 1a): We use Cartesian coordinates rather than spherical. The f-plane approximation uses a fixed value of the Coriolis parameter $f(=2\varOmega\sin\theta)$ set to correspond to mid-latitudes ($\theta=\theta_{0}$). Here $\varOmega$ is the background rotation rate.
@@
  • Geostrophic balance: The Rossby number of the flow is small in the stellar radiative zone i.e. $Ro\sim\Delta\varOmega/\varOmega\sim 0.1$ (Garaud and Arreguin, 2009). Here, $\Delta\varOmega$ is the deviation of the latitudinal differential rotation from the background $\varOmega$. Inertial forces are therefore dominated by rotation. Note that other forcings like gravitational contraction are stronger sources of $\Delta\varOmega$ (Gouhier et al., 2021) leading to larger Rossby numbers, but we do not consider them here.
@@ 2.2 The steady-state streamfunction
The independent variables representing the vertical and the meridional coordinates are pressure $p$ and distance $y$, respectively. Time is denoted by $t$. The zonal momentum, thermal wind, mass continuity and thermodynamic energy equations for a stratified hydrostatic atmosphere with momentum forcing (Showman et al., 2006) are given by, respectively,
@@
where $u$, $v$, and $w(\equiv\partial p/\partial t)$ are the zonal, meridional, and vertical wind speeds in the rotating frame of reference, $f(=2\varOmega\sin\theta)$ is the Coriolis frequency parameter, $\theta$ is the latitude and $R$ is the ideal gas constant. The momentum forcing relaxes $u$ towards $u_{F}$, and the thermal forcing relaxes $T$ towards $T^{\prime}_{E}$. The corresponding frictional drag ($\tau_{F}$) and radiative ($\tau_{R}$) relaxation times approximate the usual operators for viscous and heat diffusion. The Brunt frequency $N$ for the stratified radiative zone with a temperature distribution $T=T_{0}(p)+T^{\prime}$, where $T^{\prime}$ is a small perturbation from the reference state $T_{0}$, is given by
@@
where $C_{P}$ is the specific heat of the constituent gas at constant pressure or along isobars, $H=RT/g$ is the isothermal scale height, and $g$ is the acceleration due to gravity. The nonlinear advection terms in the momentum equation can be ignored in first approximation owing to the low Rossby number $Ro$ of the flow resulting from geostrophic balance (Vallis, 2017). We thus have four equations (1)-(4) and four variables $u$, $v$, $w$ and $T^{\prime}$.
In this work, we consider thermal forcing to be zero i.e. $T^{\prime}_{E}=0$ (Busse, 1981) to focus on the role of momentum forcing for meridional circulation. On eliminating $u$ and $T^{\prime}$ from (1)-(4) and substituting $v$ and $w$ with their corresponding expressions in terms of the meridional streamfunction $\psi$ ($v=-\partial\psi/\partial p$ and $w=\partial\psi/\partial y$), in steady-state we obtain
@@
The domain of interest is divided into two parts following the setup of Showman et al. (2006): the top layer corresponds to the near-convection zone where momentum forcing is applied (RHS of equation 1), while the deeper, stably stratified layer below is passive and only driven by the circulation in the top forced layer. The forcing velocity $u_{F}$, relaxation timescales $\tau_{R}$ and $\tau_{F}$, and the Rossby deformation radius $L_{D}(=NH/f)$ are piecewise constant along the vertical coordinate $p$, demarcating the two layers described. It is to be noted that the full two-layer setup will not be entirely solved in this work. In the following sections, the role of the upper layer is to simply inform the boundary condition on the bottom layer. However, the functional form of $u_{F}$ is mentioned here for completeness.
$u_{F}(p,y)=\begin{cases}U_{F}\cos(ly),\text{if}\quad p\leq p_{rcb}\text{ i.e. in top layer }\\ 0,\quad\quad\quad\quad\text{if}\quad p>p_{rcb}\text{ i.e. in bottom layer}\end{cases}$ (7)
where, $U_{F}$ is the velocity forcing magnitude and $p_{rcb}$ is the pressure at the radiative-convective boundary. In the passive layer, where $u_{F}=0$, equation (6) reduces to,
$L_{D}^{2}\frac{\partial^{2}\psi}{\partial y^{2}}+p^{2}\frac{\tau_{F}}{\tau_{R}}\frac{\partial^{2}\psi}{\partial p^{2}}=0.$ (8)
This equation, together with appropriate boundary conditions, gives the steady-state meridional streamfunction in a rotating hydrostatically and geostrophically balanced fluid such as the radiative zone of a Sun-like star. We see that $\psi$ admits a separable solution of the form $\psi(p,y)=\psi_{p}(p)\psi_{y}(y)$, with $\psi_{y}(y)=\cos(ly)$ i.e. periodic in latitude, where $l(=2\pi/L_{y})$ can represent the latitudinal wavenumber of the shallow jets forced in the top layer extending from the equator to the poles. Substituting the above in equation (8) we obtain an ordinary differential equation for the vertical component of the streamfunction $\psi_{p}$ as
$p^{2}\frac{\partial^{2}\psi_{p}}{\partial p^{2}}=\bar{\varOmega}^{2}\psi_{p}$ (9)
where $\bar{\varOmega}^{2}=L_{D}^{2}l^{2}\tau_{R}/\tau_{F}$. It admits the polynomial solution for $\psi$:
@@
where, $n_{1}$ and $n_{2}$ are roots of the equation $n(n-1)=\bar{\varOmega}^{2}$. Constants $A$ and $B$ may be determined from interface and boundary conditions. The vertical dependence of the steady-state streamfunction, which determines the steady-state depth of the meridional circulation, is therefore strictly dependent on $\bar{\varOmega}$ which is the product of the Burger number (Bu = $L_{D}^{2}l^{2}$) and the Prandtl number (Pr = $\tau_{R}/\tau_{F}$) as has already been established in literature e.g. Spiegel and Zahn (1992); Garaud (2002); Gilman and Miesch (2004); Garaud and Brummell (2008). For the sake of a more popularly used expression, we may rewrite $\bar{\varOmega}=(N/f)\sqrt{Pr}Hl$, explicitly showing the ratio of buoyancy and rotation frequencies. We now use this setup and formalism to explore the time-dependent behavior of the meridional circulation in a top-forced radiative zone.
@@ 3 Time-dependent zonal velocity equation
$f\frac{\partial\bar{u}}{\partial\ln p}=lR\bar{T^{\prime}}$ (13)
@@
where, $L_{D}=(NH)/f$ is the Rossby deformation radius, the log-pressure vertical coordinate $z=\ln p$ and $k=1/(L_{D}^{2}l^{2}\tau_{R})$. It can be rewritten as a linear advection-diffusion equation for the evolution of the zonal wind (RHS of equation 16). Accounting for the simplification of the radiative diffusion operator as Newtonian linear relaxation, equation (16) shares close similarity with the hyperdiffusion equation obtained for the radial profile of differential rotation by Spiegel and Zahn (1992). In its most general form, equations (16) may have $L_{D},\tau_{R},\tau_{F}$, and $k$ that dependent on $p$. However, it is possible to find useful approximate analytical solutions to equation (16) if the above parameters are constant. We apply the Dirichlet boundary condition $\bar{u}(0,t)=u_{F}$ at the top of the passive domain as an idealized form of zonal forcing. The bottom of the domain, $L$, has a free slip (no shear) boundary, and hence a Neumann boundary condition given by $u_{z}(L,t)=0$. Under these conditions, one can use solution C3 in Van Genuchten (1982) for the steady-state,
@@
$\bar{u}(z,t)=\frac{u_{F}B(z,t)}{2C(z)};m^{\pm}=\exp{\left(-\frac{k\pm j}{2k}\right)};n^{\pm}=\frac{z\pm jt}{2\sqrt{+kt}};\\ B(z,t)=m^{-}\operatorname{erfc}(n^{-})+m^{+}\operatorname{erfc}(n^{+})+\exp{\left(\frac{jL}{k}\right)}\times\\ \left[\frac{j+k}{j-k}m^{+}\operatorname{erfc}{\left(\frac{L}{\sqrt{+kt}}-n^{+}\right)}+\frac{j-k}{j+k}m^{-}\operatorname{erfc}{\left(\frac{L}{\sqrt{+kt}}-n^{-}\right)}\right]\\ -\frac{k\tau_{F}}{2}\exp{\left(-L-\frac{t}{\tau_{F}}\right)}\operatorname{erfc}{\left(\frac{2L-z-kt}{2\sqrt{+kt}}\right)}$ (18)